OpenStudy (anonymous):

Find the GCF of the monomials: 18x^2, 30x^2y, and 54y A. 18 B. 6 C. 18xy D. 6xy

OpenStudy (anonymous):

how far did you get with it?

OpenStudy (anonymous):

I don't even know where to start :/

OpenStudy (anonymous):

ok so a "greatest common factor" is the largest thing which divides all of those expressions i would suggest breaking down each monomial into its factors, then find all the shared factors, i'll break up the first one, can you try the second one? \[18x^2 = 2 \times 3 \times 3 \times x \times x\]

OpenStudy (anonymous):

have you done it?

OpenStudy (anonymous):

yes, that answer is 6

OpenStudy (anonymous):

correct :D